Derive the projection rule in a triangle ABC using cosine rule.

Grade:11

1 Answers

Vikas TU
4 years ago
Cosine Rule states that 
a^2 = b^2 + c^2 - 2bcCosA
b^2 = a^2 + c^2 - 2acCosB
c^2 = a^2 + b^2 - 2abCosC
Projecton formula gives the relatin between angles and side of a trainge 
a= b Cos C + c CosB 
B = c CosA + aCosC
c = aCosB + b CosA

Hope this helps 
Good luck